If a,b,c are the sides of the triangle ABC,such that (1+(b-c)/a)a *(1+(c-a)/b)b *(1+(a-b)/c)c>= 1


then the triangle must be


a. right angled


b. isoceles


c. obtuse angled


d.Equilateral

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Vaibhav

(1+(b-c)/a)a *(1+(c-a)/b)b *(1+(a-b)/c)c>= 1

every bracket is less than or equal to 1

but right hand side is greater than equal to 1

so its only possible when each bracket equal to 1

(1+(b-c)/a)=1

1+(b-c)/a =1

  b=c

 similerl from other brackets

 c=a

 a=b

 tiangle is equileteral

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin